# Non-Abelian Statistics for Bosonic Symmetry-Protected Topological Phases

### v3

@@ Abstract
Symmetry-protected non-Abelian (SPNA) statistics opens new frontiers in quantum statistics and enriches the schemes for topological quantum computing. In this work, we propose a novel type of SPNA statistics in one-dimensional strongly correlated bosonic symmetry-protected topological (SPT) phases and reveal its exotic universal features through a comprehensive investigation. Specifically, we show a universal result for a wide range of bosonic SPT phases described by real Hamiltonians: the SPNA statistics of topological zero modes fall into two distinct classes. The first class exhibits conventional braiding operation of hard-core bosons. Furthermore, we discover a second class of unconventional braiding statistics characterized by an exotic nonlinear transformation, featuring a fractionalization of the first class and reminiscent of the non-Abelian statistics of symmetry-protected Majorana pairs. The two distinct classes of statistics have a topological origin in classifying non-Abelian Berry phases for braiding processes of real-Hamiltonian systems, distinguished by whether the holonomy involves a reflection operation. To illustrate, we focus on a specific bosonic SPT phase with particle-hole symmetry, and demonstrate that both classes of braiding statistics can be feasibly realized in a tri-junction with or without the aid of a controlled defect. The topological zero modes are protected by unitary symmetries and are therefore immune to dynamical symmetry breaking. Numerical results support our theoretical predictions. We demonstrate how to encode logical qubits and implement both single- and two-qubit gates using the two classes of SPNA statistics. Finally, we propose feasible experimental schemes to realize these SPNA statistics and identify the parameter regimes that ensure high-fidelity braiding results, paving the way for experimental validation of our predictions and their application in quantum information science.
@@ I Introduction
The pursuit of non-Abelian statistics has been a central theme in quantum physics for decades, driven by both theoretical interest in fundamental physics and potential applications in topological quantum computing Nayak et al. (2008); Nayak and Wilczek (1996); Ivanov (2001a); Sarma et al. (2005); Alicea et al. (2011). Non-Abelian statistics revolutionizes the traditional particle classification and defines the concept of non-Abelian anyons. Among the candidates for non-Abelian anyons, Majorana zero modes (MZMs) in topological superconductors have been highly anticipated Kitaev (2001); Ivanov (2001a); Alicea et al. (2011); Alicea (2012); Aasen et al. (2016). Despite extensive experimental efforts Mourik et al. (2012); Deng et al. (2012); Rokhinson et al. (2012); Das et al. (2012); Wang et al. (2012); Churchill et al. (2013); Xu et al. (2014); Nadj-Perge et al. (2014); Chang et al. (2015); Albrecht et al. (2016); Wiedenmann et al. (2016); Bocquillon et al. (2017); Zhang et al. (2018); Wang et al. (2018); Fornieri et al. (2019); Ren et al. (2019); Jäck et al. (2019); Aghaee et al. (2023), the experimental signatures of MZMs remain illusive. Key challenges include their susceptibility to disorder Yu et al. (2021); Sarma and Pan (2021); Ahn et al. (2021); Pan and Sarma (2021) and the soft-gap problem from the proximity effect Takei et al. (2013); Stanescu et al. (2014); Liu et al. (2017). Meanwhile, although digital simulations of non-Abelian braiding have been actively studied on quantum computing platforms recently Google Quantum AI (2023); Xu et al. (2023); Iqbal et al. (2024), only analog simulations can implement non-Abelian quasiparticles and their braiding in quantum many-body phases with gap protection, which remains a challenge in experiment. These difficulties have motivated the search for alternative approaches.
@@
Our main results are as follows. We propose a new type of SPNA statistics for topological zero modes in 1D bosonic SPT phases given by a spin-exchange model, which can be mapped to the hard-core bosonic Su-Schrieffer-Heeger (SSH) model Su et al. (1979). The ground states of this model correspond to the bosonic SPT phase classified by group cohomology $H^{2}(U(1)\times\mathbb{Z}_{2}^{T},U_{T}(1))$ Chen et al. (2013). We show that the topological zero modes generically exhibit two distinct classes of SPNA statistics, which are valid for all bosonic SPT phases with real Hamiltonians. The first class is conventional braiding statistics, similar to the one of hard-core bosons. Further, we discover a second class of more unconventional braiding statistics, which is characterized by an exotic nonlinear transformation, featuring a fractionalization of the former class, and manifests emergent non-Abelian Majorana pairs. The two classes of braiding statistics can be feasibly realized using tri-junction configurations and, in particular, the fractionalized unconventional braiding is achieved with the aid of a controlled defect. The conventional and unconventional braiding statistics belong to distinct topological classes with different nontrivial non-Abelian Berry phases. Unlike in the chain geometry, the model in a tri-junction geometry is generically not exactly solvable due to strong correlation. In the tri-junction case, we numerically confirm our analytical predictions. As expected, we show that the braiding statistics is robust against symmetry-preserving disorders and does not necessitate fine-tuning of parameters. Additionally, the presence of two classes of braiding schemes can facilitate the implementation of the single- and two-qubit topological gates. Finally, we propose feasible experimental scheme to observe the SPNA statistics in the bosonic SPT phase, paving a way for experimental realization of non-Abelian quasiparticles and their braiding statistics in the near future.
@@
The remainder of this paper is organized as follows. Section II discusses the symmetry protection of topological zero modes within the bosonic SPT phase of a topological spin-exchange model. In Sec. III, we develop a generic effective Hamiltonian theory for the SPNA statistics of hard-core bosonic topological zero modes, protected by particle number conservation and particle-hole symmetry which are unitary, with the universal feature of two classes of SPNA statistics being obtained. Section IV presents a numerical verification of our prediction. The effects of dynamical symmetry breaking and the mechanism of symmetry protection for SPNA statistics are discussed in more detail in Sec. V. Section VI proposes experimental schemes for realizing the predicted SPNA statistics and highlights their high feasibility for analog quantum simulations. Finally, Sec. VII provides our conclusion and an outlook on future important issues.
@@
Figure 1: Illustration of the topological spin-exchange model and braiding schemes. Solid line between sites represents the strong spin-exchange coupling $v>0$, while the dashed line denote weak couplings $v^{\prime}>0$ with $v^{\prime}<v$. (a) The topological spin-exchange model ($N=5$) with two sublattices indicated by green and blue strips. The green and blue circles represent the left and right zero modes, respectively, localized at the ends of the system. (b) Low-energy spectrum at half filling (i.e., half of the spins are in the spin-up state) with symmetry protection (left) and symmetry breaking (right). The spectrum is ploted with coupling strengths set to $v=1$ and $v^{\prime}=0.5$. The half-filled model with symmetry protection exhibit two-fold ground-state degeneracy with an exponentially decaying energy difference $\Delta E$ between the two ground states and a finite bulk gap $\delta E$ between the ground state and the first excited state. (c) The first type of braiding scheme realizes the conventional braiding of zero modes. Green and blue arrows indicate the motion orders and directions of zero modes for braiding. The integers $l_{v}$ and $l_{h}$ represent the number of sites in the vertical direction and the half number of sites in the horizontal direction of the tri-junction (with only left-right symmetric tri-junctions considered). The whole system is half-filled. (d) The second type of braiding scheme with a local defect (encircled by the red dotted line) realizes the exotic unconventional braiding. The local defect is empty, while the rest of the tri-junction is half-filled. (e) Schematic of zero mode movement. The red site marks the central position of the localized zero mode, which shifts by two sites when coupling strengths $u$ and $v$ are adiabatically tuned. The time $T_{0}$ represents the duration of the movement. (f) The energy gap $\delta E$ (in units of $v$) between ground states and excited states of the tri-junction in the first braiding scheme as a function of the vertical ($l_{v}$) and horizontal ($l_{h}$) sizes. The orange line corresponds to varying $l_{h}$ with fixed $l_{v}=2$, while the blue dashed line corresponds to varying $l_{v}$ with fixed $l_{h}=3$. The horizontal axis denotes $l_{h}$ for the orange line and $l_{v}$ for the blue dashed line.
@@ II Topological bosonic zero modes with symmetry protection
We start in this section to investigate the symmetry protection mechanism of topological zero modes and braiding statistics in the bosonic SPT phase in both the static and dynamical regimes. We focus on a specific model with particle number conservation and particle-hole symmetry. In the static regime, symmetries of the microscopic Hamiltonian guarantee the presence of zero modes in the topological phase. We also examine symmetry protection during the dynamical evolution of zero modes. This is crucial for understanding SPNA statistics, as braiding is an inherently dynamical process. For the SPNA statistics to be well-defined, symmetry protection must be maintained throughout the evolution and be consistent with the dynamical evolution of zero modes. We show that unitary symmetries provide this necessary protection, fulfilling both requirements.
II.1 Zero modes in the topological spin-exchange model
@@
where $v^{\prime}$ and $v$ are non-negative real parameters. This model can be mapped to the hard-core bosonic Su-Schrieffer-Heeger (SSH) model via a Holstein-Primakoff transformation. In this mapping, the spin up (down) state is identified with the bosonic vacuum state $\left|{0}\right\rangle$ (occupied state $\left|{1}\right\rangle=b^{\dagger}\left|{0}\right\rangle$), where the hard-core bosonic operators are defined as $b^{\dagger}_{i}=\sigma_{i}^{-}\equiv(\sigma^{x}_{i}-i\sigma^{y}_{i})/2$ and $b_{i}=\sigma_{i}^{+}\equiv(\sigma_{i}^{x}+i\sigma_{i}^{y})/2$. The mapped Hamiltonian is then given by
@@
In the topological phase ($v>v^{\prime}$), the model hosts two zero modes, each localized at one end of the chain. This feature is transparent in the dimerized limit, where one of $v$ or $v^{\prime}$ vanishes while the other remains finite. In this limit, pairs of the bulk spins form singlet states, i.e., dimers, while two unpaired zero modes localize at the ends. The ground states then factorize into a product of decoupled bulk dimers and edge modes:
$\left|{\Psi_{A}}\right\rangle=\frac{(b^{\dagger}_{1,a})^{n_{1}}(b^{\dagger}_{N,b})^{n_{2}}}{2^{N/2}}\prod_{j=1}^{N-1}\left(b^{\dagger}_{j,b}-b^{\dagger}_{j+1,a}\right)\left|{\Uparrow}\right\rangle,$ (3)
where $\left|{\Uparrow}\right\rangle=\left|{\uparrow}\right\rangle^{\otimes 2N}$ denotes the vacuum state with all sites unoccupied. The zero modes can be either empty ($n_{i}=0$) or occupied ($n_{i}=1$), resulting in a four-fold degenerate ground-state subspace. In the Fock space representation, the ground-state basis is $\left|{n_{1},n_{2}}\right\rangle$}. In contrast, the trivial phase ($v<v^{\prime}$) has a unique ground state in the dimerized limit:
$\left|{\Psi_{B}}\right\rangle=\frac{1}{2^{N/2}}\prod_{j}\left(b^{\dagger}_{j,a}-b^{\dagger}_{j,b}\right)\left|{\Uparrow}\right\rangle.$ (4)
We formally label the creation operators of the left and right zero modes as $\beta_{1}^{\dagger}$ and $\beta_{2}^{\dagger}$, which in the dimerized limit reduce to $\beta_{1}^{\dagger}=b^{\dagger}_{1,a}$ and $\beta_{2}^{\dagger}=b^{\dagger}_{N,b}$. Away from the dimerized limit, the zero-mode creation operators can no longer be expressed as single local boson operators. Instead, each mode has support over an entire sublattice and decays exponentially from one end of the chain. In the case of $v>>v^{\prime}$ and the thermodynamic limit $N\gg 1$, the zero-mode operators read:
@@
where $\mathcal{N}\approx\sqrt{1-\left(\frac{v^{\prime}}{v}\right)^{2}}$ is the normalization factor. It is obvious that zero-mode operators obey the same commutation relations as hard-core bosons. When the system size is finite, the finite-size effect hybridizes the left and right zero modes, opening an exponentially small energy splitting $\Delta E\propto\left(\frac{v^{\prime}}{v}\right)^{N}$ [left panel of Fig. 1(b)], which is negligible in practice for typical large system size. The zero end modes are then protected by the bulk topological gap $\delta E$.
@@ II.2 Symmetry protection of zero modes
$U_{\theta}=\text{exp}(i\theta\sum_{i}b_{i}^{\dagger}b_{i}),\quad C=\prod_{i}(b_{i}+b_{i}^{\dagger}),$ (7)
where the subscrip $i$ collectively denotes all relevant indices for simplicity. The two unitary symmetries are sufficient to protect the zero modes. To demonstrate this, we employ the effective-Hamiltonian approach to symmetry analysis: after integrating out the local bosonic modes by restricting to the ground-state subspace, the residual dynamics is governed by symmetry-allowed terms constructed from the zero-mode operators $\beta_{1}$, $\beta_{2}$, $\beta_{1}^{\dagger}$, and $\beta_{2}^{\dagger}$. The most general terms in the effective Hamiltonian capable of lifting the zero modes take the form
@@
Symmetry protection manifests in the fact that these perturbations fail to commute with the symmetry operators, as can be directly verified
$C\epsilon_{j}(\beta_{j}^{\dagger}\beta_{j}-1/2)C^{-1}$ $=$ $-\epsilon_{j}(\beta_{j}^{\dagger}\beta_{j}-1/2),$
$U_{\theta}(m_{j}\beta_{j}+m_{j}^{*}\beta_{j}^{\dagger})U_{\theta}^{-1}$ $=$ $e^{-i\theta}m_{j}\beta_{j}+e^{i\theta}m_{j}^{*}\beta_{j}^{\dagger},$
$U_{\theta}(\Delta\beta_{1}\beta_{2}+\Delta^{*}\beta_{1}^{\dagger}\beta_{2}^{\dagger})U_{\theta}^{-1}$ $=$ $e^{-2i\theta}\Delta\beta_{1}\beta_{2}+e^{2i\theta}\Delta^{*}\beta_{1}^{\dagger}\beta_{2}^{\dagger}.$
To illustrate the necessity of both unitary symmetries for protecting the zero modes and the resulting ground-state degeneracy, we numerically compute the low-energy spectrum at half filling in the presence of a symmetry-breaking term,
@@
where $\delta_{1}$ and $\delta_{2}$ are real numbers. This Hamiltonian does not commute with $C$ when $\sin\omega t\neq 0$ but commute with $C^{A}$ and $U_{\theta}$ at arbitrary time $t$. When $\omega\gg\delta_{1(2)}$, the effective Hamiltonian obtained by Floquet theory (Appendix A) reads
@@
where the second term destroys the zero modes. Therefore, although the anti-unitary symmetry can protect zero modes in the static regime, their properties are lost during time evolution. In Sec. V, we further study dynamical symmetry breaking caused by complex random hopping disorders.
The role of anti-unitary symmetries in this context is to enable a more refined mathematical classification. Specifically, without the complex conjugation $K$, the symmetry operators $U_{\theta}$ and $C$ do not commute, yielding the symmetry group $U(1)\rtimes\mathbb{Z}_{2}$ with group cohomology classification $H^{2}(U(1)\rtimes\mathbb{Z}_{2},U(1))=\mathbb{Z}_{2}$. Including $K$ allows one to choose a representation $\tilde{U}_{\theta}=\text{exp}(i\theta\sum_{i}(b^{\dagger}_{i}b_{i}-\frac{1}{2}))$ such that $\tilde{U}_{\theta},C^{A}]=0$. In this case, the symmetry group becomes $U(1)\times\mathbb{Z}^{T}_{2}$, and the corresponding classification is $H^{2}(U(1)\times\mathbb{Z}^{T}_{2},U_{T}(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Chen et al. (2012). We stress that, in this example, anti-unitary symmetry is not necessary for the physical protection of zero modes. Even with only unitary symmetries, the topological classification remains non-trivial. Consequently, we focus primarily on unitary-symmetry protection, but will revisit anti-unitary symmetries in Sec. V, where dynamical symmetry breaking is discussed. The unitary-symmetry protection renders the zero modes robust against dynamical symmetry breaking and is essential for well-defined braiding results.

### v4

@@ Abstract
Symmetry-protected non-Abelian (SPNA) statistics opens new frontiers in quantum statistics and enriches the schemes for topological quantum computing. In this work, we propose a new paradigm of SPNA statistics in one-dimensional correlated bosonic symmetry-protected topological (SPT) phases and uncover exotic universal features from a systematic investigation. In particular, we show that for generic bosonic SPT phases described by real Hamiltonians, the SPNA statistics of topological zero modes fall into two distinct classes. The first class exhibits conventional braiding of hard-core bosonic zero modes. Furthermore, we discover a second class of unconventional braiding statistics characterized by a nonlinear transformation, featuring a fractionalization of the first class and reminiscent of the non-Abelian statistics of symmetry-protected Majorana pairs. The two distinct classes of statistics have topological origin in classifying non-Abelian Berry phases for braiding processes of real-Hamiltonian systems, distinguished by whether the holonomy involves a reflection operation. To illustrate, we focus on a specific bosonic SPT phase with particle-hole symmetry, and demonstrate that both classes of braiding statistics can be feasibly realized in a tri-junction with and without the aid of a controlled defect, respectively. Analytic and numerical results are given. We demonstrate how to encode logical qubits and implement both single- and two-qubit gates using the two classes of SPNA statistics. Finally, we propose feasible experimental schemes to observe these predictions and identify the parameter regimes for the high-fidelity braiding, paving the way for the experimental observation of our results in the near future.
@@ I Introduction
The pursuit of non-Abelian statistics has been a central theme in quantum physics for decades, driven by both theoretical interest in fundamental physics and potential applications in topological quantum computing Wilczek (1990); Moore and Read (1991); Nayak and Wilczek (1996); Ivanov (2001a); Kitaev (2003); Sarma et al. (2005); Nayak et al. (2008); Alicea et al. (2011). Non-Abelian statistics revolutionizes the traditional particle classification and defines the concept of non-Abelian anyons. Among the candidates for non-Abelian anyons, Majorana zero modes (MZMs) in topological superconductors have been highly anticipated Read and Green (2000); Kitaev (2001); Ivanov (2001a); Alicea et al. (2011); Alicea (2012); Sarma et al. (2015); Aasen et al. (2016); Sato and Fujimoto (2016). Despite extensive experimental efforts Mourik et al. (2012); Deng et al. (2012); Rokhinson et al. (2012); Das et al. (2012); Wang et al. (2012); Churchill et al. (2013); Xu et al. (2014); Nadj-Perge et al. (2014); Chang et al. (2015); Albrecht et al. (2016); Wiedenmann et al. (2016); Bocquillon et al. (2017); Zhang et al. (2018); Wang et al. (2018); Fornieri et al. (2019); Ren et al. (2019); Jäck et al. (2019); Aghaee et al. (2023), the experimental signatures of MZMs remain illusive. Key challenges include their susceptibility to disorder Yu et al. (2021); Sarma and Pan (2021); Ahn et al. (2021); Pan and Sarma (2021) and the soft-gap problem from the proximity effect Takei et al. (2013); Stanescu et al. (2014); Liu et al. (2017). Meanwhile, although digital simulations of non-Abelian braiding have been actively studied on quantum computing platforms recently Google Quantum AI (2023); Xu et al. (2023); Iqbal et al. (2024), only analog simulations can implement non-Abelian quasiparticles and their braiding in quantum many-body phases with gap protection, which remains a challenge in experiment. These difficulties have motivated the search for alternative approaches.
@@
Our core results are as follows. We predict that the topological zero modes in the 1D bosonic SPT phase exhibit two fundamental classes of SPNA statistics – a universal feature of the bosonic SPT phases with real Hamiltonians and in sharp contrast to the free-fermion topological phases. The first class is the conventional braiding statistics, similar to the one of hard-core bosons. Remarkably, we discover a second class of unconventional braiding statistics, which is characterized by an exotic nonlinear transformation, featuring a fractionalization of the former class and a reminiscent of braiding two Majorana pairs. The conventional and unconventional braiding statistics belong to distinct topological classes with different nontrivial non-Abelian Berry phases, with which the non-commuting braiding matrices can be implemented on a single qubit given by two bosonic zero modes. The predicted exotic braiding statistics unveil an intrinsic dynamical feature of bosonic SPT phases, providing new insights into their characterization and fundamental nature. Further, we consider a topological spin-exchange model, which can be mapped to the hard-core bosonic Su-Schrieffer-Heeger (SSH) model Su et al. (1979), to illustrate the general theory. In particular, the two classes of SPNA statistics are shown to be achievable by braiding the zero modes via a tri-junction with and without the assistance of a local defect put in the junction point, respectively. This provides a highly feasible way toward the experimental realization of SPNA statistics. Unlike a single chain model, the model in a tri-junction geometry is generically not exactly solvable due to strong correlation, and our analytical predictions are confirmed with numerical results. As expected, we show that the braiding statistics is protected and robust against local symmetry preserving disorders. Additionally, the presence of two classes of non-Abelian statistics facilitates the implementation of the single- and two-qubit topological gates. Finally, we propose real experimental setting to observe the SPNA statistics in the bosonic SPT phase, paving a way for experimental realization in the near future.
@@
The remainder of this paper is organized as follows. Section II introduces the topological zero modes in the bosonic SPT phase of a topological spin-exchange model, with the symmetry protection of the topological phase being studied in both static and dynamical regimes. In Sec. III, we develop a generic effective Hamiltonian theory for the SPNA statistics of the bosonic zero modes, protected by particle number conservation and particle-hole symmetry which are unitary, with the universal feature of two classes of SPNA statistics being obtained. Section IV presents a numerical verification of our prediction. The effects of dynamical symmetry breaking and the mechanism of symmetry protection for SPNA statistics are discussed in more detail in Sec. V. Section VI proposes experimental schemes for realizing the predicted SPNA statistics and highlights their high feasibility for analog quantum simulations. Finally, the conclusion and outlook for future issues are given in Sec. VII.
@@
Figure 1: Illustration of the topological spin-exchange model and braiding schemes. Solid line between sites represents the strong spin-exchange coupling $v>0$, while the dashed line denote weak couplings $v^{\prime}>0$ with $v^{\prime}<v$. (a) The topological spin-exchange model ($N=5$) with two sublattices indicated by green and blue strips. The green and blue circles represent the left and right zero modes, respectively, localized at the ends of the system. (b) Low-energy spectrum at half filling (i.e., half of the spins are in the spin-up state) with symmetry protection (left) and symmetry breaking (right). The spectrum is ploted with coupling strengths set to $v=1$ and $v^{\prime}=0.5$. The half-filled model with symmetry protection exhibit two-fold ground-state degeneracy with an exponentially decaying energy difference $\Delta E$ between the two ground states and a finite bulk gap $\delta E$ between the ground state and the first excited state. (c) The first type of braiding scheme realizes the conventional braiding of zero modes. Green and blue arrows indicate the motion orders and directions of zero modes for braiding. The integers $l_{v}$ and $l_{h}$ represent the number of sites in the vertical direction and the half number of sites in the horizontal direction of the tri-junction (with only left-right symmetric tri-junctions considered). The whole system is half-filled. (d) The second type of braiding scheme with a local defect (encircled by the red dotted line) realizes the exotic unconventional braiding. The local defect is empty, while the rest of the tri-junction is half-filled. (e) Schematic of zero mode movement. The red site marks the position of the localized zero mode, which shifts by two sites when coupling strengths $u$ and $v$ are adiabatically tuned. The time $T_{0}$ represents the duration of the movement. (f) The energy gap $\delta E$ (in units of $v$) between ground states and excited states of the tri-junction in the first braiding scheme as a function of the vertical ($l_{v}$) and horizontal ($l_{h}$) sizes. The orange line corresponds to varying $l_{h}$ with fixed $l_{v}=2$, while the blue dashed line corresponds to varying $l_{v}$ with fixed $l_{h}=3$. The horizontal axis denotes $l_{h}$ for the orange line and $l_{v}$ for the blue dashed line.
@@ II Topological bosonic zero modes with symmetry protection
We start in this section to investigate the symmetry protection mechanism of topological zero modes and braiding statistics in the bosonic SPT phase in both the static and dynamical regimes. For clarity, we focus on a spin-exchange model with alternating coupling strengths, whose symmetries include unitary particle-hole symmetry, particle-number conservation, and an anti-unitary symmetry. Although both unitary and antiunitary symmetries can protect the topology of bosonic SPT phases and ensure the existence of zero modes in the static setting, we show that only unitary symmetries can protect the topological phases in the dynamical regime. Since the braiding is inherently dynamical, the SPNA statistics are well defined when the protecting symmetry is preserved throughout the entire braiding operation, both for the static topological phase and for the dynamical braiding evolution, a condition that we show is generically satisfied by unitary symmetries.
II.1 Bosonic zero modes in the topological spin-exchange model
@@
where $v^{\prime}$ and $v$ are non-negative real parameters. This model can be mapped to the hard-core bosonic Su-Schrieffer-Heeger (SSH) model Su et al. (1979) via Holstein-Primakoff transformation. In this mapping, the spin up (down) state is identified with the bosonic vacuum state $\left|{0}\right\rangle$ (occupied state $\left|{1}\right\rangle=b^{\dagger}\left|{0}\right\rangle$), where the hard-core bosonic operators are defined as $b^{\dagger}_{i}=\sigma_{i}^{-}\equiv(\sigma^{x}_{i}-i\sigma^{y}_{i})/2$ and $b_{i}=\sigma_{i}^{+}\equiv(\sigma_{i}^{x}+i\sigma_{i}^{y})/2$. The mapped Hamiltonian is then given by
@@
In the topological phase ($v>v^{\prime}$), the model hosts two hard-core bosonic zero modes, each localized at one end of the chain. We formally label the creation operators of the left and right zero modes as $\beta_{1}^{\dagger}$ and $\beta_{2}^{\dagger}$. The zero modes can be either empty ($n_{i}=0$) or occupied ($n_{i}=1$), resulting in a four-fold degenerate ground-state subspace spanned by the states $\lvert n_{1},n_{2}\rangle$}. In the case of $v>>v^{\prime}$ and the thermodynamic limit $N\gg 1$, the zero-mode operators read:
@@
where $\mathcal{N}\approx\sqrt{1-\left(\frac{v^{\prime}}{v}\right)^{2}}$ is the normalization factor. The zero-mode operators obey the same commutation relations as hard-core bosons, commute with $H_{0}$ within the ground-state manifold, and reduce to $\beta_{1}^{\dagger}=b^{\dagger}_{1,a}$ and $\beta_{2}^{\dagger}=b^{\dagger}_{N,b}$ in the dimerized limit. When the system is finite, the finite-size effect hybridizes the left and right zero modes, opening an exponentially small energy splitting $\Delta E\propto\left(\frac{v^{\prime}}{v}\right)^{N}$ [left panel of Fig. 1(b)], which is negligible for typical large system size. The zero end modes are then protected by the bulk topological gap $\delta E$.
@@ II.2 Symmetry protection of zero modes
$U_{\theta}=\text{exp}(i\theta\sum_{i}b_{i}^{\dagger}b_{i}),\quad C=\prod_{i}(b_{i}+b_{i}^{\dagger}).$ (5)
The two unitary symmetries are sufficient to protect the zero modes. To demonstrate this, we employ the effective-Hamiltonian approach to symmetry analysis: the dynamics of zero modes is governed by symmetry-allowed terms constructed from the zero-mode operators $\beta_{1}$, $\beta_{2}$, $\beta_{1}^{\dagger}$, and $\beta_{2}^{\dagger}$. The most general terms in the effective Hamiltonian capable of lifting the zero modes take the form
@@
Symmetry protection is manifested in the fact that these perturbations do not commute with the symmetry operators. To illustrate the necessity of both unitary symmetries for protecting the zero modes and the resulting ground-state degeneracy, we numerically compute the low-energy spectrum at half filling in the presence of a symmetry-breaking term,
@@
where $\delta_{1,2}$ are real numbers. This Hamiltonian commutes with $C^{A}$ and $U_{\theta}$ at arbitrary time $t$ but breaks $C$ for $\sin\omega t\neq 0$. When $\omega\gg\delta_{1(2)}$, the effective Hamiltonian obtained by Floquet theory (Methods VIII.1) reads
@@
The second term in $H_{\text{eff}}$ destroys the zero modes, implying that the anti-unitary symmetry in the static regime is generically broken down in the dynamical regime. In Sec. V, we further study dynamical symmetry breaking caused by complex random hopping disorders.
The role of anti-unitary symmetries in this context is to enable a more refined mathematical classification. Specifically, without the complex conjugation $K$, the symmetry operators $U_{\theta}$ and $C$ do not commute, yielding the symmetry group $U(1)\rtimes\mathbb{Z}_{2}$ with group cohomology classification $H^{2}(U(1)\rtimes\mathbb{Z}_{2},U(1))=\mathbb{Z}_{2}$. Including $K$ allows one to choose a representation $\tilde{U}_{\theta}=\text{exp}(i\theta\sum_{i}(b^{\dagger}_{i}b_{i}-\frac{1}{2}))$ such that $\tilde{U}_{\theta},C^{A}]=0$. In this case, the symmetry group becomes $U(1)\times\mathbb{Z}^{T}_{2}$, and the corresponding classification is $H^{2}(U(1)\times\mathbb{Z}^{T}_{2},U_{T}(1))=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ Chen et al. (2012, 2013). We stress that, in this example, anti-unitary symmetry is not necessary for the physical protection of zero modes. Even with only unitary symmetries, the topological classification remains non-trivial. Consequently, we focus primarily on unitary-symmetry protection, but will revisit anti-unitary symmetries in Sec. V, where dynamical symmetry breaking is discussed. The unitary-symmetry protection renders the zero modes robust against dynamical symmetry breaking and is essential for well-defined braiding results.